(1 point) Rock band The Rolling Stones have played scores of concerts in the last twenty years. For 30 randomly selected Rolling Stones concerts, the mean gross earnings is 2.65 million dollars. Part a) Assuming a population standard deviation gross earnings of 0.53 million dollars, obtain a 99% confidence interval for the mean gross earnings of all Rolling Stones concerts (in millions). Please carry at least three decimal places in intermediate steps. Give your answer to the nearest 3 decimal places. Confidence interval: ( , ).
Answer #1

Confidence Interval :-
X̅ ± Z( α /2) σ / √ ( n )
Z(α/2) = Z (0.01 /2) = 2.576
2.65 ± Z (0.01/2 ) * 0.53/√(30)
Lower Limit = 2.65 - Z(0.01/2) 0.53/√(30)
Lower Limit = 2.401
Upper Limit = 2.65 + Z(0.01/2) 0.53/√(30)
Upper Limit = 2.899
99% Confidence interval is ( 2.401 , 2.899 )